OpenStudy (anonymous):

Hey, I am graphing a Sin graph and I just want some help with the Phase Shift section. The initial graph problem is 2sin(4x+pi/2)-3. I believe the formula for the Phase Shift is -c/b so I did -pi/2/4. This gives me a compound fraction and I am confused as to where to go from here.

jimthompson5910 (jim_thompson5910):

phase shift = -c/b phase shift = -(pi/2)/(4) phase shift = -(pi/2)/(4/1) phase shift = -(pi/2)*(1/4) phase shift = -(pi*1)/(2*4) phase shift = -pi/8

OpenStudy (anonymous):

thank you for the help!